A consumer has $25 in budget to purchase goods X & Y. Assume Px = 3 and Py=1. She has a coupon that allows her to buy-one-get-one-free for up to 5 units of X. What is the maximum amount of X that she can buy?


1
Expert's answer
2020-02-11T09:25:32-0500

budget =$25

Px= 3

Py= 1

budget line, "bl=Px\/Py" = 3/1= 3

If the consumer buys five units of X he gets five more units, therefore, the amount of X he can buy is given by:

"3*5units + (15)" =30 units